If one were to to apply Bohr Model to a particle of mass “m” and charge “q” moving in a plane under the influence of a magnetic field “B”, the energy of charge particle in the nth level will be.??

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
m *v*r = n*h/2\pi
&
q *v*B = m*v²/r
So
q*B = m*v/r
q*B(nh/2\pim*v) = m*v
(½ *m* v²) = nhqB/4\pim
E = n (h*q*B/4* \large \pi *m)
 
Regards
Arun (askIITians forum expert)